OpenStudy (anonymous):

write two related division facts for each multiplication fact. (-3)*(+6)=+18 . (-5)*(+4)=-20 . (+7)*(-3)=-21 . (+8)*(+6)=+48.

OpenStudy (anonymous):

Look at your first fact.. there is something wrong with one of the signs.

OpenStudy (anonymous):

2nd fact -20/-5 = 4 and -20/4 = -5 3rd fact -21/7 = -3 and -21/-3 = 7 4th fact 48/8 = 6 and 48/6 = 8

OpenStudy (anonymous):

oh ya it was -6 not +6

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

18/-6 = -3 and 18/-3 = -6

OpenStudy (anonymous):

thanks